OpenStudy (anonymous):

A rectangular dog pen has an area of 108 square feet. The width of the pen is 9 feet. What is the length?

OpenStudy (anonymous):

i got 6. i did 9*9=18 18/108=6 9*6=54 54+54=108

OpenStudy (anonymous):

A=area w=width l=length A=w × l plug in the given values... 108=9 × l now divide both sides by 9, and you get.... 12= l so the missing length is 12.

OpenStudy (anonymous):

student_basil is right. As for GinjaTheNinja, the length would have had to be more than the width, so 6 doesn't make sense.

OpenStudy (anonymous):

ohh thanks